# Odd fake $\mathbb{Q}$-homology quadrics exist

### v1

@@ 2. Definitions and basic properties
It will be called an odd homology quadric if instead the intersection form is odd (hence diagonalizable with diagonal entries $(+1,-1)$).
@@
An even homology quadric $S$ must be minimal, since it contains no $(-1)$-curves. Hence, by surface classification, for an even homology quadric
@@
$\mathfrak{A}_{5}$ $\langle 60,5\rangle$ $[5,5,5]$ $[2,2,2,3]$ 1 $(\mathbb{Z}_{10})^{2}$
$\mathfrak{A}_{5}$ $\langle 60,5\rangle$ $[3,3,5]$ $[2,2,2,2,2]$ 2 $(\mathbb{Z}_{2})^{3}\times\mathbb{Z}_{6}$
$\mathfrak{S}_{4}\times\mathbb{Z}_{2}$ $\langle 48,48\rangle$ $[2,4,6]$ $[2,2,2,2,2,2]$ 3 $(\mathbb{Z}_{2})^{4}\times\mathbb{Z}_{4}$
G(32) $\langle 32,27\rangle$ $[2,2,4,4]$ $[2,2,2,4]$ 2 $(\mathbb{Z}_{2})^{2}\times\mathbb{Z}_{4}\times\mathbb{Z}_{8}$
@@
$D_{4}\times\mathbb{Z}_{2}$ $\langle 16,11\rangle$ $[2,2,2,4]$ $[2,2,2,2,2,2]$ 4 $(\mathbb{Z}_{2})^{3}\times(\mathbb{Z}_{4})^{2}$
@@
$(\mathbb{Z}_{2})^{3}$ $\langle 8,5\rangle$ $[2,2,2,2,2]$ $[2,2,2,2,2,2]$ 5 $(\mathbb{Z}_{2})^{4}\times(\mathbb{Z}_{4})^{2}$
@@
Then $[K_{S}]\in Num(S)$ is not 2-divisible if and only if $\Phi_{1}$ is not 2-divisible.
@@ 3.1. Final remark
There remains the question whether there is a $G$- invariant Thetacharacteristic $\theta$ inside $Pic^{3}(C)$. We have seen that it must be $H^{0}(\theta)=0$, and that $\theta\otimes L_{j}$ are $G$ invariant, for $j=1,2$.

### v2

@@ Acknowledgement.
Thanks to Kyoung-Seog Lee for asking a question concerning the proof, which lead me to add a few lines of explanation before Proposition 3.1.
@@ 2. Definitions and basic properties
It will be called an odd homology quadric if instead the intersection form is odd (hence diagonalizable with diagonal entries (+1,1)(+1,-1)).
@@
An even homology quadric $S$ must be minimal, since it contains no (1)(-1)-curves. Hence, by surface classification, for an even homology quadric
@@
$\mathfrak{A}_{5}$ $\langle 60,5\rangle$ $[5,5,5]$ $[2,2,2,3]$ ? 1 $(\mathbb{Z}_{10})^{2}$
$\mathfrak{A}_{5}$ $\langle 60,5\rangle$ $[3,3,5]$ $[2,2,2,2,2]$ ? 2 $(\mathbb{Z}_{2})^{3}\times\mathbb{Z}_{6}$
$\mathfrak{S}_{4}\times\mathbb{Z}_{2}$ $\langle 48,48\rangle$ $[2,4,6]$ $[2,2,2,2,2,2]$ ? 3 $(\mathbb{Z}_{2})^{4}\times\mathbb{Z}_{4}$
G(32) $\langle 32,27\rangle$ $[2,2,4,4]$ $[2,2,2,4]$ ? 2 $(\mathbb{Z}_{2})^{2}\times\mathbb{Z}_{4}\times\mathbb{Z}_{8}$
@@
$D_{4}\times\mathbb{Z}_{2}$ $\langle 16,11\rangle$ $[2,2,2,4]$ $[2,2,2,2,2,2]$ ? 4 $(\mathbb{Z}_{2})^{3}\times(\mathbb{Z}_{4})^{2}$
@@
$(\mathbb{Z}_{2})^{3}$ $\langle 8,5\rangle$ $[2,2,2,2,2]$ $[2,2,2,2,2,2]$ ? 5 $(\mathbb{Z}_{2})^{4}\times(\mathbb{Z}_{4})^{2}$
@@
Then $K_{S}]\in Num(S)$ is not 2-divisible if and only if $\Phi_{1}$ is not 2-divisible.
@@
In fact, $\Phi_{1}$ is 2-divisible in $Pic(S)$ if and only if on $C_{1}\times C_{2}$ there is a $G$-linearized square root $\mathcal{L}$} of the line bundle $p_{1}^{*}(K_{C_{1}})$.
$\mathcal{L}$} would be of the form $p_{1}^{*}(\theta)+p_{2}^{*}(M)$, where $M$ is a 2-torsion line bundle on $C_{2}$. Since $M$ is a torsion line bundle, it is a pull back form $Jac(C_{2})$, and as such it admits a $G$-linearization for the induced action of $G$ on $Jac(C_{2})$. Hence the existence of a $G$-linearization on $\mathcal{L}$} is equivalent to the existence of a $G$-linearization on $\theta$.
@@ 3.1. Final remark
There remains the question whether there is a $G$- invariant Thetacharacteristic $\theta$ inside $Pic^{3}(C)$. We have seen that it must be $H^{0}(\theta)=0$, and that $\theta\otimes L_{j}$ are $G$ - invariant, for $j=1,2$.